L^{A}TEX News

Issue 10, December 1998

Five years of L^{A}TEX 2ε

Since this is the 10th edition of L^{A}TEX News, the (no longer) New Standard L^{A}TEX must have hit the streets almost this long ago. In fact it was only the beta-version that some people got just in time for Christmas 1993, and since then there has been a lot of tidying-up and smoothing of rough edges (not to mention a few bug fixes!).

Maybe it is time for something more radically different to emerge and be hungrily adopted by the world; but don’t panic, we shall be maintaining what you have now for a long time yet. Amongst the more polite things that have been written about our efforts, we found that this quote (somewhat censored to protect the guilty) well reflects some of our feelings about working on L^{A}TEX over the years: the mere existence of L^{A}TEX 2ε is a great miracle.

Restructuring the L^{A}TEX distribution

Since the (once) ‘new’ standard L^{A}TEX has reached such a venerable age, we are reviewing the way in which the system is presented to the world.

An early intention is to define, given the wide variety of good packages now available, what now constitutes a useful installation of L^{A}TEX. We also hope that such a definition will help document portability if it leads to a future in which a L^{A}TEX class designer can reasonably assume that a known list of facilities will be there for all users (so that each class need not supply them).

As a first small step towards this definition, we shall replace the latex/packages subdirectory on CTAN. This directory was a curious mixture of the important, such as the L^{A}TEX tools, that any self-respecting L^{A}TEX installation ought to have, and the esoteric or experimental.

The esoterica from packages will be moved to new locations, as follows:

    expl3 to latex/exptl/project
    mfnfss to latex/contrib/supported/mfnfss

The subdirectory that replaces packages will be called latex/required; all the other sub-directories of packages will be moved there.

L^{A}TEX Project on the Internet

A new latex-project.org domain has been registered. The web site is not yet fully functional but the old L^{A}TEX pages from CTAN are available at

                                                   http://www.latex-project.org/ and the L^{A}TEX bug

reporting address has been changed to

                                                   latex-bugs@latex-project.org.

Restructuring the L^{A}TEX package licenses

Several people have requested an easy mechanism for the distribution of L^{A}TEX packages and other software “under the same conditions as L^{A}TEX”. The old legal.txt file was unsuitable as a general licence as it referred to specific L^{A}TEX authors, and to specific files. Therefore, in this release legal.txt contains just the

copyright notice and a reference to the new L^{A}TEX Project Public License (LPPL) for the distribution and modification conditions. The tools, graphics, and mfnfss packages also now refer to this license in their distribution notices.

Support for Cyrillic encodings

Basic Cyrillic support, as announced in L^{A}TEX News 9, is now finally an official part of L^{A}TEX. It includes support for the following standard Cyrillic font encodings (this list may grow): T2A T2B T2C X2.

It also includes various Cyrillic input encodings (20 in total, including commonly used variants and Mongolian Cyrillic encodings). This provides platform independent and sophisticated basic support for high-quality typesetting in various Cyrillic-based languages.

For further information see the file cyrguide.tex.

Tools distribution

The varioref package has been extended to support textual page references to a range of objects: e.g., if eq-first and eq-last are the label names for the first and last equation in a sequence, then you can now write

                                                     see~\vrefrange{eq-first}{eq-last}

This results in different text depending on whether both labels fall on the same page.

Some additional user commands, as well as building-blocks for writing private extensions, are described in the accompanying documentation.

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 1998, all rights reserved.